OpenStudy (anonymous):

Precalc help please! Prove 1 + sec^(2)x sin^(2)x = sec^(2)x

hartnn (hartnn):

do you know this: \(\sin^2x+\cos^2 x=1\) ?

OpenStudy (anonymous):

yes that's the pythagorean identity

hartnn (hartnn):

correct, now try to isolate sin^2 x from that (subtract cos^2 x from both sides) what do u get ?

OpenStudy (anonymous):

1 - cos^2x

hartnn (hartnn):

correct, now plug that in your left side equation what u get?

OpenStudy (anonymous):

1+sec^2 x (1-cos^2 x)

hartnn (hartnn):

good, distribute sec^2 x \(1+\sec^2 x -\sec^2 x \cos^2 x\) got this ?

hartnn (hartnn):

can you simplify that further ?

OpenStudy (anonymous):

how did we distribute sec2x?

hartnn (hartnn):

\(\large ☻(☺+♥) = ☻☺+☻♥ \\ \text {this is distributing}\)

OpenStudy (anonymous):

wait i got it!

hartnn (hartnn):

\(\sec^2 x (1-cos^2 x ) = \sec^2 x - \sec^2 x \cos^2 x\) good :)

hartnn (hartnn):

can you tell me what sec^2 x cos^2 x = ... ?

OpenStudy (anonymous):

this is where i get lost :(

hartnn (hartnn):

did you know \(\Large \sec x =\dfrac{1}{\cos x}\)

OpenStudy (anonymous):

yes and cos x = 1/ sec x

hartnn (hartnn):

which gives us \(\Large \sec x \cos x = 1\) isn't it ?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

i think i got it!

OpenStudy (anonymous):

1 + sec2x - 1 1 - 1 = 0 so, sec2x = sec2x

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

AAHHHH! I can't believe I did it! Thank you soooooo much!

hartnn (hartnn):

and since you are new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

hartnn (hartnn):

you're most welcome ^_^

OpenStudy (anonymous):

Can you check my work on one more problem? I feel like the process was too short to be right

hartnn (hartnn):

sure

OpenStudy (anonymous):

-tan^2 x + sec^2 x = 1 -tan^2 x + 1 + tan^2 x the - tan and tan would cancel out so, 1 = 1

hartnn (hartnn):

you wanted to prove -tan^2 x + sec^2 x = 1 ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

ok, what you did is correct, but you have used sec^2 x = 1+ tan^2 x

hartnn (hartnn):

if you are in doubts, don't use that

hartnn (hartnn):

use these : tan x = sin x/ cos x sec x = 1/cos x

hartnn (hartnn):

then try to simplify by taking the common denominator

OpenStudy (anonymous):

so since tan is negative would it be - sinx/cosx or would i use the even odd identity

hartnn (hartnn):

- tan x = -sin x/ cos x -tan^2 x = - sin^2 x/ cos^2 x

OpenStudy (anonymous):

alright that makes sense

OpenStudy (anonymous):

so is this right so far, - sin^2x/cos^2x + 1/cos^2x 1- sin^2x

hartnn (hartnn):

denominator cos^2 x is common so you combine the numerator \(\Large \dfrac{1-\sin^2 x}{\cos^2x}\)

OpenStudy (anonymous):

okay

hartnn (hartnn):

and whats 1-sin^2 x = .. .?

OpenStudy (anonymous):

cos^2x

hartnn (hartnn):

good, so both numerator and denominator = cos^2 x which gives you 1 :)

OpenStudy (anonymous):

you're a genius!

hartnn (hartnn):

not actually, but thanks for the compliment :)

OpenStudy (anonymous):

anytime! :)